# Palm distributions of superposed point processes for statistical inference

### v1

@@ Acknowledgments
$M^{(k)}_{\Phi}(B_{1},\ldots,B_{k})$ $=\sum_{q=1}^{k}\sum_{\{J_{1},\ldots,J_{q}\}}\prod_{h=1}^{q}\int_{\times_{m\in J_{h}}B_{m}}\eta(\bm{x}_{J_{h}})\;\prod_{m\in J_{h}}\mathrm{d}x_{m}$
$=\sum_{q=1}^{k}\sum_{\{J_{1},\ldots,J_{q}\}}\int_{B_{1}\times\cdots\times B_{k}}\prod_{h=1}^{q}\eta(\bm{x}_{J_{h}})\;\mathrm{d}x_{1}\,\ldots\,\mathrm{d}x_{k}$
$=\int_{B_{1}\times\cdots\times B_{k}}\sum_{q=1}^{k}\sum_{\{J_{1},\ldots,J_{q}\}}\prod_{h=1}^{q}\eta(\bm{x}_{J_{h}})\;\mathrm{d}x_{1}\,\ldots\,\mathrm{d}x_{k},$
and the thesis follows, with the introduction of the allocation variables $\bm{t}\in\mathcal{T}_{k}$.
Proof of Theorem S2
The formulation of the sncp as cluster process in (5) yields a clustering structure at the latent level. Based on this, the generating mechanism of the points can be described in terms of the clustering structure and the cluster-specific parameters of the kernel $\kappa$. Specifically, consider the event $((\theta^{*}_{1},N^{*}_{1}),\ldots,(\theta^{*}_{C},N^{*}_{C}))$, where $C$ denotes the number of clusters among the $N$ points of $\Phi$, $\theta^{*}_{h}$ represents the parameter of kernel $\kappa$ in cluster $h$, and $N^{*}_{h}$ is the number of points of $\Phi$ in cluster $h$, given a uniform random ordering of the clusters. It holds that $N=\sum_{h=1}^{C}N^{*}_{h}$ and the cluster-specific parameters $\theta^{*}_{h}$’s are distinct. Denote with $\bm{\theta}^{*}=(\theta^{*}_{1},\ldots,\theta^{*}_{C})$. The main computation concerns the probability of the event just described, that is
$\mathsf{P}((\theta^{*}_{1},N^{*}_{1}),\ldots,(\theta^{*}_{C},N^{*}_{C}))$ $=\mathsf{E}\left[\mathsf{P}((\theta^{*}_{1},N^{*}_{1}),\ldots,(\theta^{*}_{C},N^{*}_{C})\,|\,\Lambda)\right]$
$=\frac{1}{C!}\mathsf{E}\left[\int_{(\mathbb{X}\times\mathbb{R}_{+})^{C}}\prod_{h=1}^{C}\frac{1}{{N^{*}_{h}}!}e^{-\gamma_{h}}\gamma_{h}^{N^{*}_{h}}\delta_{\theta^{*}_{h}}(\theta_{h})\cdot\prod_{j>C}e^{-\gamma_{j}}\Lambda^{C}(\mathrm{d}\bm{\theta}\,\mathrm{d}\bm{\gamma})\right]$
$=\frac{1}{C!}\mathsf{E}\left[\int_{(\mathbb{X}\times\mathbb{R}_{+})^{C}}e^{-\int_{\mathbb{X}\times\mathbb{R}_{+}}s\Lambda(\mathrm{d}x\,\mathrm{d}s)}\cdot\prod_{h=1}^{C}\frac{1}{{N^{*}_{h}}!}\gamma_{h}^{N^{*}_{h}}\delta_{\theta^{*}_{h}}(\theta_{h})\Lambda^{C}(\mathrm{d}\bm{\theta}\,\mathrm{d}\bm{\gamma})\right],$
@@
$\{(\bm{\theta},\bm{\gamma})\in(\mathbb{X}\times\mathbb{R}_{+})^{C}:\;\theta_{i}=\theta_{j}\text{ for }i\neq j\},$
@@
$\mathsf{P}((\theta^{*}_{1},N^{*}_{1}),\ldots,(\theta^{*}_{C},N^{*}_{C}))$
$\quad=\frac{1}{C!}\mathsf{E}\left[\int_{(\mathbb{X}\times\mathbb{R}_{+})^{C}}e^{-\int_{\mathbb{X}\times\mathbb{R}_{+}}s\Lambda(\mathrm{d}x\,\mathrm{d}s)}\cdot\prod_{h=1}^{C}\frac{1}{{N^{*}_{h}}!}\gamma_{h}^{N^{*}_{h}}\delta_{\theta^{*}_{h}}(\theta_{h})\Lambda^{(C)}(\mathrm{d}\bm{\theta}\,\mathrm{d}\bm{\gamma})\right]$
$\quad=\frac{1}{C!}\int_{(\mathbb{X}\times\mathbb{R}_{+})^{C}}\mathsf{E}\left[e^{-\int_{\mathbb{X}\times\mathbb{R}_{+}}s\Lambda^{!}_{\bm{\theta},\bm{\gamma}}(\mathrm{d}x\,\mathrm{d}s)}\right]e^{-\sum_{h=1}^{C}\gamma_{h}}\prod_{h=1}^{C}\frac{1}{{N^{*}_{h}}!}\gamma_{h}^{N^{*}_{h}}\delta_{\theta^{*}_{h}}(\theta_{h})\;M_{\Lambda}^{(C)}(\mathrm{d}\bm{\theta}\,\mathrm{d}\bm{\gamma})$
$\quad=\frac{1}{C!}\int_{\mathbb{R}_{+}^{C}}\mathsf{E}\left[e^{-\int_{\mathbb{X}\times\mathbb{R}_{+}}s\Lambda^{!}_{\bm{\theta}^{*},\bm{\gamma}}(\mathrm{d}x\,\mathrm{d}s)}\right]\prod_{h=1}^{C}\frac{1}{{N^{*}_{h}}!}e^{-\gamma_{h}}\gamma_{h}^{N^{*}_{h}}\;M_{\Lambda}^{(C)}(\mathrm{d}\bm{\theta}^{*}\,\mathrm{d}\bm{\gamma}).$ (S19)
@@
$\mathsf{P}((\theta^{*}_{1},N^{*}_{1}),\ldots,(\theta^{*}_{C},N^{*}_{C}))$ $=\frac{1}{C!}\mathsf{E}\left[e^{-\int_{\mathbb{X}\times\mathbb{R}_{+}}s\Lambda(\mathrm{d}x\,\mathrm{d}s)}\right]\int_{\mathbb{R}_{+}^{C}}\prod_{h=1}^{C}\frac{1}{{N^{*}_{h}}!}e^{-\gamma_{h}}\gamma_{h}^{N^{*}_{h}}\;\nu^{C}(\mathrm{d}\bm{\theta}^{*}\,\mathrm{d}\bm{\gamma})$
@@
$\mathsf{P}((\theta^{*}_{1},\mathcal{C}_{1}),\ldots,(\theta^{*}_{C},\mathcal{C}_{C}))=\binom{N}{N^{*}_{1},\ldots,N^{*}_{k}}^{-1}\mathsf{P}((\theta^{*}_{1},N^{*}_{1}),\ldots,(\theta^{*}_{C},N^{*}_{C})).$
@@
$\mathsf{P}(\mathcal{C}_{1},\ldots,\mathcal{C}_{C})=\frac{1}{N!}\frac{1}{C!}e^{-\int_{\mathbb{X}\times\mathbb{R}_{+}}(1-e^{-\gamma})\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)}\prod_{h=1}^{C}\int_{\mathbb{X}\times\mathbb{R}_{+}}e^{-\gamma_{h}}\gamma_{h}^{N^{*}_{h}}\;\nu(\mathrm{d}\theta_{h}\,\mathrm{d}\gamma_{h}).$
Observe that $\mathsf{P}(\{\mathcal{C}_{1},\ldots,\mathcal{C}_{C}\})$ is obtained by multiplying by $C!$ the previous expression. In the statement of the theorem, we stress that this probability distribution describes the number of points $N$ as well.

### v2

@@ Acknowledgments
$M^{(k)}_{\Phi}(B_{1},\ldots,B_{k})$ $=\sum_{q=1}^{k}\sum_{\{J_{1},\ldots,J_{q}\}}\prod_{h=1}^{q}\int_{\times_{m\in J_{h}}B_{m}}\eta(\underline{x}_{J_{h}})\;\prod_{m\in J_{h}}\mathrm{d}x_{m}$
$=\sum_{q=1}^{k}\sum_{\{J_{1},\ldots,J_{q}\}}\int_{B_{1}\times\cdots\times B_{k}}\prod_{h=1}^{q}\eta(\underline{x}_{J_{h}})\;\mathrm{d}x_{1}\,\ldots\,\mathrm{d}x_{k}$
$=\int_{B_{1}\times\cdots\times B_{k}}\sum_{q=1}^{k}\sum_{\{J_{1},\ldots,J_{q}\}}\prod_{h=1}^{q}\eta(\underline{x}_{J_{h}})\;\mathrm{d}x_{1}\,\ldots\,\mathrm{d}x_{k},$
and the thesis follows, with the introduction of the allocation variables $\underline{t}\in\mathcal{T}_{k}$.
Proof of Theorem˜S2
The formulation of the sncp as cluster process in (3) yields a clustering structure at the latent level. Based on this, the generating mechanism of the points can be described in terms of the clustering structure and the cluster-specific parameters of the kernel $\kappa$. Specifically, consider the event $((\theta^{*}_{1},N^{*}_{1}),\ldots,(\theta^{*}_{C},N^{*}_{C}))$, where $C$ denotes the number of clusters among the $N$ points of $\Phi$, $\theta^{*}_{h}$ represents the parameter of kernel $\kappa$ in cluster $h$, and $N^{*}_{h}$ is the number of points of $\Phi$ in cluster $h$, given a uniform random ordering of the clusters. It holds that $N=\sum_{h=1}^{C}N^{*}_{h}$ and the cluster-specific parameters $\theta^{*}_{h}$’s are distinct. Denote with $\underline{\theta}^{*}=(\theta^{*}_{1},\ldots,\theta^{*}_{C})$. The main computation concerns the probability of the event just described, that is
${\rm pr}((\theta^{*}_{1},N^{*}_{1}),\ldots,(\theta^{*}_{C},N^{*}_{C}))$ $=E\left[{\rm pr}((\theta^{*}_{1},N^{*}_{1}),\ldots,(\theta^{*}_{C},N^{*}_{C})\,|\,\Lambda)\right]$
$=\frac{1}{C!}E\left[\int_{(\mathbb{X}\times\mathbb{R}_{+})^{C}}\prod_{h=1}^{C}\frac{1}{{N^{*}_{h}}!}e^{-\gamma_{h}}\gamma_{h}^{N^{*}_{h}}\delta_{\theta^{*}_{h}}(\theta_{h})\cdot\prod_{j>C}e^{-\gamma_{j}}\Lambda^{C}(\mathrm{d}\underline{\theta}\,\mathrm{d}\underline{\gamma})\right]$
$=\frac{1}{C!}E\left[\int_{(\mathbb{X}\times\mathbb{R}_{+})^{C}}e^{-\int_{\mathbb{X}\times\mathbb{R}_{+}}s\Lambda(\mathrm{d}x\,\mathrm{d}s)}\cdot\prod_{h=1}^{C}\frac{1}{{N^{*}_{h}}!}\gamma_{h}^{N^{*}_{h}}\delta_{\theta^{*}_{h}}(\theta_{h})\Lambda^{C}(\mathrm{d}\underline{\theta}\,\mathrm{d}\underline{\gamma})\right],$
@@
$\{(\underline{\theta},\underline{\gamma})\in(\mathbb{X}\times\mathbb{R}_{+})^{C}:\;\theta_{i}=\theta_{j}\text{ for }i\neq j\},$
@@
${\rm pr}((\theta^{*}_{1},N^{*}_{1}),\ldots,(\theta^{*}_{C},N^{*}_{C}))$
$\quad=\frac{1}{C!}E\left[\int_{(\mathbb{X}\times\mathbb{R}_{+})^{C}}e^{-\int_{\mathbb{X}\times\mathbb{R}_{+}}s\Lambda(\mathrm{d}x\,\mathrm{d}s)}\cdot\prod_{h=1}^{C}\frac{1}{{N^{*}_{h}}!}\gamma_{h}^{N^{*}_{h}}\delta_{\theta^{*}_{h}}(\theta_{h})\Lambda^{(C)}(\mathrm{d}\underline{\theta}\,\mathrm{d}\underline{\gamma})\right]$
$\quad=\frac{1}{C!}\int_{(\mathbb{X}\times\mathbb{R}_{+})^{C}}E\left[e^{-\int_{\mathbb{X}\times\mathbb{R}_{+}}s\Lambda^{!}_{\underline{\theta},\underline{\gamma}}(\mathrm{d}x\,\mathrm{d}s)}\right]e^{-\sum_{h=1}^{C}\gamma_{h}}\prod_{h=1}^{C}\frac{1}{{N^{*}_{h}}!}\gamma_{h}^{N^{*}_{h}}\delta_{\theta^{*}_{h}}(\theta_{h})\;M_{\Lambda}^{(C)}(\mathrm{d}\underline{\theta}\,\mathrm{d}\underline{\gamma})$
$\quad=\frac{1}{C!}\int_{\mathbb{R}_{+}^{C}}E\left[e^{-\int_{\mathbb{X}\times\mathbb{R}_{+}}s\Lambda^{!}_{\underline{\theta}^{*},\underline{\gamma}}(\mathrm{d}x\,\mathrm{d}s)}\right]\prod_{h=1}^{C}\frac{1}{{N^{*}_{h}}!}e^{-\gamma_{h}}\gamma_{h}^{N^{*}_{h}}\;M_{\Lambda}^{(C)}(\mathrm{d}\underline{\theta}^{*}\,\mathrm{d}\underline{\gamma}).$ (S37)
@@
${\rm pr}((\theta^{*}_{1},N^{*}_{1}),\ldots,(\theta^{*}_{C},N^{*}_{C}))$ $=\frac{1}{C!}E\left[e^{-\int_{\mathbb{X}\times\mathbb{R}_{+}}s\Lambda(\mathrm{d}x\,\mathrm{d}s)}\right]\int_{\mathbb{R}_{+}^{C}}\prod_{h=1}^{C}\frac{1}{{N^{*}_{h}}!}e^{-\gamma_{h}}\gamma_{h}^{N^{*}_{h}}\;\nu^{C}(\mathrm{d}\underline{\theta}^{*}\,\mathrm{d}\underline{\gamma})$
@@
${\rm pr}((\theta^{*}_{1},\mathcal{C}_{1}),\ldots,(\theta^{*}_{C},\mathcal{C}_{C}))=\binom{N}{N^{*}_{1},\ldots,N^{*}_{k}}^{-1}{\rm pr}((\theta^{*}_{1},N^{*}_{1}),\ldots,(\theta^{*}_{C},N^{*}_{C})).$
@@
${\rm pr}(\mathcal{C}_{1},\ldots,\mathcal{C}_{C})=\frac{1}{N!}\frac{1}{C!}e^{-\int_{\mathbb{X}\times\mathbb{R}_{+}}(1-e^{-\gamma})\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)}\prod_{h=1}^{C}\int_{\mathbb{X}\times\mathbb{R}_{+}}e^{-\gamma_{h}}\gamma_{h}^{N^{*}_{h}}\;\nu(\mathrm{d}\theta_{h}\,\mathrm{d}\gamma_{h}).$
Observe that ${\rm pr}(\{\mathcal{C}_{1},\ldots,\mathcal{C}_{C}\})$ is obtained by multiplying by $C!$ the previous expression. In the statement of the theorem, we stress that this probability distribution describes the number of points $N$ as well.